if a+b+c=0 then find the value of a2+b2+c2/c2-ab.
Please tell the answer

Grade:8

1 Answers

Arun
4 years ago
Dear student
 
a + b + c = 0
 
a + b = – c
 
hence (a +b)^2 =c^2
 
hence a^2 + b^2 = c^2 – 2ab
 
now coming to question
 
(c^2 – 2ab + c^2 )/ c^2 – ab = 2(c^2 – ab) /(c^2 – ab) = 2